OpenStudy (anonymous):

Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 2x - y = 5 x + 2y = 15

OpenStudy (anonymous):

Do you need help explaining it or solving it?

OpenStudy (anonymous):

solving pleeeeeasseeee

OpenStudy (anonymous):

Okay so you need to choose a variable and equation in which to substitute. Let's take equation two and sub x in.

OpenStudy (anonymous):

1 Rearrange equation 2. x+2y=15 x=15-2y

OpenStudy (anonymous):

2. Now we sub x into equation 1. 2(15-2y)-y=5 30-4y-y=5 -5y=5-30 -5y=-25 y=5

OpenStudy (anonymous):

3. Now that we have solved for y we can find x. x=15-2*5 =15-10 =5 So x and y are both 5